OpenStudy (anonymous):

Solve for x.. x^2+3x-40=0

OpenStudy (king):

x=5,-8

OpenStudy (anonymous):

Can you explain please

OpenStudy (hoblos):

x^2+3x-40=(x-5)(x+8)=0 => x=5 or x=-8

OpenStudy (king):

yeah sure see use quadratic equation alpha/beta are roots the equation is alpha/beta=(-b+or -sqrt [b^2 -4ac])/2a a,b,c are coefficents ax+by+c=0